OpenStudy (anonymous):

derivative of y= the square root of x(3x-1)

OpenStudy (anonymous):

\[y=\sqrt{x}(3x-1)\]

OpenStudy (dumbcow):

you have to use product rule

OpenStudy (dumbcow):

(fg)' = f'g + fg' let f = sqrt(x) = x^(1/2) g = 3x-1 f' = 1/2x^(-1/2) g' = 3

jimthompson5910 (jim_thompson5910):

y = sqrt(x)*(3x-1) y' = d/dx(sqrt(x)*(3x-1)) y' = d/dx(sqrt(x))*(3x-1) + sqrt(x)*d/dx(3x-1) y' = d/dx(x^(1/2))*(3x-1) + sqrt(x)*d/dx(3x-1) y' = (1/2)x^(1/2-1)*(3x-1) + sqrt(x)*(3) y' = (1/2)x^(-1/2)*(3x-1) + 3*sqrt(x) y' = (1/(2x^(1/2)))*(3x-1) + 3*sqrt(x) y' = (1/(2*sqrt(x)))*(3x-1) + 3*sqrt(x) So \[\Large y^{\prime} = \frac{1}{2\sqrt{x}}(3x-1) + 3\sqrt{x}\]